Determine whether the two line are parallel, perpendicular or neither.
𝑥 − 3𝑦 = 6
𝑦 = 1 x+4
___
3 ­


Answer by Boreal(15235)

look at them in slope intercept form
the first is 3y=x-6
y=(1/3)x-2
second is y=x+4
look at slopes, which are (1/3) and 1
they aren't the same and their product is not -1. They are neither parallel nor perpendicular.